# LLM3: Large Language Model-based Task and Motion Planning with Motion Failure Reasoning

### v2

@@ V Conclusions
In this paper, we introduced LLM3, a new TAMP framework powered by the Large Language Model (LLM). Unlike conventional TAMP frameworks, which typically entail significant efforts to explicitly design domain-specific symbolic task planners and manually interface them with motion planners, LLM3 leverages the rich knowledge encoded in and the powerful reasoning capability processed by LLMs to (i) propose action sequences based on task goals without requiring a prior planning domain, (ii) automatically generate continuous action parameters for the robot motion planner, and more importantly, (iii) refine task and/or motion plans in response to motion planning failures.
Following validation through various simulations and experiments, we demonstrated that LLM3 effectively produced and refined task and motion plans for box-packing problems, exhibiting promising potential in addressing previously unspecified tasks. Our study also revealed that although the LLM can generate action parameters more efficiently than random samplers, it still necessitated multiple feedback iterations and motion planner calls. Looking ahead, the incorporation of in-context learning or fine-tuning techniques holds promise for further enhancing its efficiency, representing a crucial step towards empowering robots to tackle emerging tasks in real-world scenarios.
Acknowledgement: We thank Hang Li, Lecheng Ruan, Junyi Liang for their fruitful discussions and support. This work was supported in part by the National Natural Science Foundation of China (No.52305007).

### v3

@@ V Conclusions
In this paper, we introduced LLM3, a new TAMP framework powered by the pre-trained LLM. LLM3 leverages the rich knowledge encoded in and the powerful reasoning capability processed by LLMs to (i) propose action sequences based without requiring a prior planning domain, (ii) generate continuous action parameters for the robot motion planner, and more importantly, (iii) refine task and/or motion plans in response to motion planning failures.
Following validation through various simulations and experiments, we demonstrated that LLM3 effectively produced and refined task and motion plans for box-packing problems, exhibiting promising potential in addressing previously unspecified tasks. Our study also revealed that although the pre-trained LLM can generate action parameters more efficiently than random samplers, it still necessitated multiple feedback iterations and motion planner calls. Looking ahead, the incorporation of in-context learning or fine-tuning techniques holds promise for further enhancing its efficiency, representing a crucial step towards empowering robots to tackle emerging tasks in real-world scenarios.